OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The second step would be to: Select one: a. Divide both sides by 3 to simplify. b. Re-write the equation using the definition of logarithms. c. Add (x – 7) to both sides to simplify. d. Either A or C

OpenStudy (luigi0210):

Do you know your log rules for solving?

OpenStudy (solomonzelman):

I am doing this.... \(\normalsize\color{black}{ \log_2x + \log_2(x – 7) = 3}\) \(\normalsize\color{black}{ \log_2x + \log_2(x – 7) = 3\log_22}\) \(\normalsize\color{black}{ \log_2x + \log_2(x – 7) = \log_22^3}\) \(\normalsize\color{black}{ \log_2x + \log_2(x – 7) = \log_29}\) \(\normalsize\color{black}{ \log_2x(x – 7) = \log_29}\) \(\normalsize\color{black}{ \log_2x^2-7x= \log_29}\) \(\normalsize\color{black}{ x^2-7x=9}\) .....

OpenStudy (solomonzelman):

convert to base 2, and equalize.

OpenStudy (solomonzelman):

I hate this type of questions "what would be your second step and why" man... just post a solution :P

OpenStudy (anonymous):

how do you get \[\3log_{3}\]

OpenStudy (solomonzelman):

I multiplied by log(2)2, because log(number)number =1 for every number, besides when number=1

OpenStudy (solomonzelman):

and not \(\normalsize\color{red}{ \log_3}\) but \(\normalsize\color{black}{ \log_2}\)

OpenStudy (anonymous):

i thought i put 2 sorry

OpenStudy (solomonzelman):

Yeah I realized.

OpenStudy (solomonzelman):

I can't say what the second step is, it depends on the first step, doesn't it ?

OpenStudy (anonymous):

yea it depends what you do first i would think it is D

OpenStudy (solomonzelman):

It is not, and none of the options that it includes for the first step :)

OpenStudy (solomonzelman):

your first step is ??

OpenStudy (solomonzelman):

Hint: ` log(a)+log(b) = log ( a × b ) `

OpenStudy (anonymous):

take the log of each side?\

OpenStudy (solomonzelman):

you already have the side in logs .... I gave you everything to figure it out.